# SUREL+: Moving from Walks to Sets for Scalable Subgraph-based Graph Representation Learning

### 17 Dec 2023 v2

@@ 3.4. Set Neural Encoders
After joining sampled node sets for each query $Q$, the resulting $(\mathcal{S}_{Q},\mathcal{Z}_{Q})$ acts as the query-induced subgraph $\mathcal{G}_{Q}$ and then is fed into neural encoders for prediction. The mini-batch training procedure of multiple queries is summarized in Algorithm 1. Next, we introduce neural encoders supported by SUREL+.
@@
Here, $enc(\cdot)$ encodes query-level structural features $\mathcal{Z}_{Q,x}$ using a multi-linear perception. If node attributes are present, they can be appended to the structural features as $\mathcal{Z}_{Q,u}||X_{u}$. AGGR is used to aggregate the encoded features, which can be any neural encoders applied on sets such as mean/sum/max pooling or set transformers. Currently, SUREL+ supports the implementations of AGGR in mean pooling, LSTM (Hamilton et al., 2017a), and attention (Veličković et al., 2018). Note that the LSTM applies random permutations to the elements in the set before encoding them as a sequence; while the attention first computes soft attention scores based on the output of $enc(\cdot)$ for each set element and then performs attention-score-weighted pooling. Sec. 4.4 empirically demonstrates that the choice of AGGR has non-trivial effects on prediction performance. Lastly, a fully connected layer takes the readout $h_{Q}$ as input to make the final prediction $\hat{y}_{Q}$. In our experiments, all SGRL tasks are formulated as binary classification, and thus Binary Cross Entropy is used as the loss function $\mathcal{L}$.
@@
3       Generate negative training queries (if not given) $\{\bar{Q}_{B}\}$ by random sampling and put them into $\mathcal{Q}_{B}$;
@@ 4. Evaluation
  • Regarding space and time complexity, how much improvement can SUREL+ achieve by adopting sets instead of walks compared to the SOTA SGRL framework SUREL?
  • Can SUREL+ provide comparable prediction performance to all baselines using or not using SGRL methods?
  • How sensitive is SUREL+ to different choices of set samplers, structure encoders, and set neural encoders?
  • How do sparse storage SpG and parallelism in SpJoin operator benefit the overall performance of SUREL+?
@@ 4.1. Experiment Setup
Extensive experiments have been performed to evaluate SUREL+ using homogeneous, heterogeneous, and higher-order homogeneous graphs on three types of tasks: link prediction, relation type prediction, and higher-order pattern prediction. A homogeneous graph is a graph that does not include node/link types, while a heterogeneous graph includes various node/link types. Higher-order graphs are hypergraphs in our setting that contain hyperedges connecting two or more nodes.
Datasets Table 2 summarizes the statistics of datasets used to benchmark SGRL methods. Five datasets are selected from the Open Graph Benchmark (OGB, (Hu et al., 2020)) for link and relation type prediction, including social networks of citation - citation2 and collaboration - collab; biological network of protein interaction - ppa and blood vessels - vessel; and one heterogeneous academic network ogb-mag, which contains node types of paper (P), author (A) and their extracted relations. The vessel dataset is a large ($>$3M nodes), sparse, biological graph recently constructed from mouse brains (Paetzold et al., 2021) and has unique significance for examining GRL in scientific discovery. The structure of vessels illustrates the spatial organization of the brain’s microvasculature, which can be used for early detection of neurological disorders, e.g. Alzheimer’s and stroke. Two hypergraph datasets collected by (Benson et al., 2018) are used for higher-order pattern prediction: DBLP-coauthor is a temporal hypergraph, where each hyperedge denotes a time-stamped paper connecting all its authors. tags-math contains sets of tags applied to questions on the website math.stackexchange.com, represented as hyperedges. For higher-order pattern prediction tasks, the number of hyperedges is the main computation bottleneck, in which one may connect more than two nodes. Two industry-level graphs, criteo-click with 16.5M records of online banner ads clicking (Diemert et al., 2017) and twitter-2010 with 1.5B following relations of users (Kwak et al., 2010) are used to examine the model scalability for real-world applications.
@@
Baselines We consider two classes of baselines. Canonical GNNs: GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017a), GraphSAINT (Zeng et al., 2020) and their variants with the prefix ‘H*’ that are directly applied for heterogeneous graphs with node types and for hypergraphs through clique expansion. R-GCN (Schlichtkrull et al., 2018) performs relational message passing on heterogeneous graphs; Scalable GNNs: and Cluster-GCN (Chiang et al., 2019); SGRL models: SEAL (Zhang and Chen, 2018; Zhang et al., 2021), GDGNN (Kong et al., 2022), and SUREL (Yin et al., 2022). SEAL adopts online subgraph sampling due to its intractable space needs for offline subgraph extraction. Fig. 3 (a) compares the time cost for subgraph sampling across different SGRL methods. We use all baselines’ official implementations with hyperparameter tuning to match their reported results.
Hyperparameters By default, SUREL+ uses the walk-based sampler, the structural encoder LP and the better set neural encoder tuned between mean pooling and attention. SUREL+ adopts a 2-layer MLP as $enc(\cdot)$ in Eq. (2) followed by a 2-layer MLP classifier to map set-aggregated representations for final predictions. Default training hyperparameters: learning rate lr=1e-3 with the early stopping of 5 epochs, dropout p=0.1, Adam (Kingma and Ba, 2015) as the optimizer. Analysis of parameters $M$ and $m$ to control the walk-based sampler and $K$ to control the metric-based sampler and selection of structure encoders and set neural encoders are studied in Sec. 4.4.
@@ 4.2. Prediction Accuracy Comparison
Table 3 shows the prediction performance of different methods. SGRL models significantly outperform canonical GNNs on these six link prediction benchmarks, especially on two challenging biological datasets ppa and vessel. Predicting links in biological datasets requires richer structural information that canonical GNNs have limited expressive power to capture. Within SGRL models, SUREL+ achieves comparable performance to SUREL and outperforms SEAL, which validates the effectiveness of the proposed set-based subgraph representation. For predictions of relation type and higher-order pattern, we observe additional performance gains (+2$\sim$13%) from SUREL+ compared to SUREL on three of the four datasets. A large performance gap exists between canonical GNNs and SGRL models, particularly in the higher-order case. This demonstrates the inherent limitations of canonical GNNs to make predictions of complex relations involving multiple nodes.
@@ 4.3. Efficiency and Scalability Analysis
Table 4 compares model runtime and memory usage on the four largest benchmarks. SUREL+ offers a reasonable training time compared with canonical GNNs. It shows clear improvement in inference compared to the current SOTA framework SUREL (3-11$\times$ speedups across all datasets) and its predecessor SEAL ($\sim$20$\times$ speedups). SUREL+ achieves comparable and even lower RAM usage than canonical GNNs. Compared to other SGRL models, it can save up to half of RAM with lower usage of GPU DRAM, since set-based subgraphs eliminate node duplicates and their associated structural features. Tables 1 and 4 analytically and empirically show that the key factor of SUREL+ scaling up to billion-size graphs is its set-based representation with the sparse design, while GCN (full adjacency matrix), SEAL (complex subgraph extraction) and SUREL (dense walks with duplicate nodes) are all out of memory (OOM) on twitter-2010.
@@
Fig. 5(a) reports the time cost of different samplers with multithreading on citation2. Fig. 5(b) shows memory consumption to store different types of sampled data (walks in SUREL (Yin et al., 2022) or sets in SUREL+) and their associated structural features (LPs, SPDs, PPR scores). Compared to SUREL sampler(Yin et al., 2022), the walk-based sampler in SUREL+ is more scalable and only adds one extra minute for encoding and converting data in SpG format (slash/dash marked in Fig. 5(a)), while achieving 6.94×6.94$\times$6.94 ×, 3.63×3.63$\times$3.63 × and 4.12×4.12$\times$4.12 × memory savings on three OGB datasets, respectively. Those memory savings are crucial for model scalability as they reduce the workload of data transfer from CPU to GPU and save GPU operations from encoding duplicated nodes, which dominates the time cost of the online stage and leads to the efficiency improvement of SUREL+ in Table 4. In addition, the metric-based sampler that adopts PPR scores has better scaling performance with more threads. When PPR scores or SPDs are adopted as structural features, SUREL+ further reduces the memory cost, though they often harm prediction performance slightly.
Note that in the above comparison of memory cost, compressed structural features are adopted both in SUREL (locally) and SUREL+ (globally), i.e., the secondary index based on SFptr $D$ and $D_{\text{SF}}$, which achieve compression of 493×493$\times$493 ×, 11318×11318$\times$11318 ×, 19527×19527$\times$19527 × on three datasets listed in Fig. 5(b), when one adopts LPs as structural features.
@@
Fig. 7 shows the speedups and throughput of the SpJoin operator for constructing query-level structural features via multithreading, where the join operation in SURE for nodes in walks (Yin et al., 2022) is used for comparison. SUREL employs a hash-based search for walk joins, which has unfavorable memory access patterns and suffers from imbalanced workloads due to inconsistent searching times across threads. SUREL+ gains more benefits from multithreading, thanks to sparse arithmetic operations and batch-wise load balancing used in SpJoin.
@@
Figure 6. Comparison of Runtime, Memory Consumption across Different Offline Processing Strategies (the walk-based sampler: $m=4,M=200$, the metric-based sampler: $K=150$). The areas highlighted break down the total consumption w.r.t. (a) sampling, structure encoding, sparse object construction; (b) structural features, node indices/pointers, and sampled walks (SUREL sampler only).
@@ 4.4. Comparison between Different Set Samplers, Structural Features and Set Neural Encoders
Table 5 shows the prediction performance and inference runtime by adopting different combinations of structure encoders and set neural encoders. Landing probabilities (LPs) as structural features perform the best on all three OGB datasets while being the slowest for inference. By recording the landing probabilities over different steps of walks, LPs provide structural information in finer granularity than scalar values of SPDs and PPR scores. The adopted link prediction task might favor more local information held by LPs and SPDs than global information carried by PPR scores. The authors conjecture that other tasks that rely on more global information may favor PPR scores. In comparison, no set neural encoder is always a winner. Attention seems to perform the best on average while slower than mean pooling. LSTM is the slowest. On the two social networks (citation2 and collab), mean pooling can provide comparable prediction results with much fewer parameters. However, prediction on the biological network (ppa) requires more expressive and complicated encoders, where LSTM and attention are favored as they can model more complex interactions between nodes in the joint set $\mathcal{S}_{Q}$.
Fig. 8 compares prediction results and inference time by using different hyperparameters $m,M$, and $K$ of set samplers, which heavily affects the coverage of sampled neighborhoods and computation overhead. The performance consistently increases if the walk-based sampler uses a larger $M$, but is not guaranteed for a larger $m$ (broader exploration). Better coverage with a larger $K$ is usually beneficial for the metric-based sampler over citation2 but not for collab, which is due to different characteristics of these two datasets and is also observed by (Yin et al., 2022). In general, small sampling parameters $m~{}(2\sim 4),M~{}(100\sim 400)$ and $K~{}(50\sim 200)$ can yield satisfactory performance with fast inference speed that achieves the trade-off between accuracy and scalability.
Table 5. Prediction Performance and Inference Runtime of SUREL+ with Different Combinations of Structure Features (LP, SPD, PPR) and Set Neural Encoders (Mean, LSTM, Attn.). The best and the second best are highlighted in bold and underlined accordingly.
@@ 5. Conclusion
In this work, we propose a novel framework SUREL+ for scalable subgraph-based graph representation learning. SUREL+ avoids costly query-specific subgraph extraction by decoupling it into sampled node sets with structural features whose join can function as query-induced subgraphs for prediction. SUREL+ benefits from the reusability and compactness of pre-sampled node sets across different queries. Compared to the SOTA SGRL framework SUREL, the set-based representation of SUREL+ substantially reduces space and time complexity by avoiding heavy node duplicates in sampled walks. To handle irregularly sized node sets, SUREL+ designs a customized sparse storage SpG and a sparse join operator SpJoin, providing memory-efficient storage with fast access. In addition, SUREL+ adopts a modular design, enabling users to choose different set samplers, structure encoders, and set neural encoders flexibly based on the nature of their SGRL tasks. Extensive experiments on three types of prediction tasks over nine real-world graph benchmarks show that SUREL+ significantly improves scalability, memory efficiency, and prediction accuracy compared to current SGRL methods and canonical GNNs.

### 27 Dec 2023 v3

@@ 3.4. Set Neural Encoders
After joining node sets for each query $Q$, the resulting $(\mathcal{S}_{Q},\mathcal{Z}_{Q})$ acts as the query-induced subgraph $\mathcal{G}_{Q}$ and then is fed into a neural encoder for prediction. The mini-batch training procedure of multiple queries is summarized in Algorithm 1. Next, we introduce neural encoders supported by SUREL+.
@@
Here, $enc(\cdot)$ encodes query-level structural features $\mathcal{Z}_{Q,x}$ using a multi-linear perception (MLP). If node attributes are present, they can be appended after structural features as $\mathcal{Z}_{Q,u}||X_{u}$. AGGR is used to aggregate the encoded features, which can be any neural encoders applicable to sets such as mean/sum/max pooling or set transformers. Currently, SUREL+ supports the implementations of AGGR in mean pooling, LSTM (Hamilton et al., 2017a), and attention (Veličković et al., 2018). Note that, the LSTM applies random permutations to the elements in the set before encoding them as a sequence; while the attention first computes soft attention scores based on the output of $enc(\cdot)$ for each set element and then performs attention-score-weighted pooling. Sec. 4.4 empirically demonstrates that the choice of AGGR has non-trivial effects on prediction performance. Lastly, a fully connected layer takes the readout $h_{Q}$ as input to make the final prediction $\hat{y}_{Q}$. In our experiments, all SGRL tasks are formulated as binary classification, and thus Binary Cross Entropy is used as the loss function $\mathcal{L}$.
@@
3       Generate negative training queries (if not given) $\{...,\bar{Q},...\}$ by random sampling and put them into $\mathcal{Q}_{B}$;
@@ 4. Evaluation
  • Regarding space and time complexity, how much improvement can SUREL+ achieve by adopting node sets instead of walks compared to the SOTA framework SUREL?
  • Can SUREL+ provide comparable prediction performance to all baselines using or not using subgraph-based methods?
  • How sensitive is SUREL+ to choices of different set samplers, structure encoders, and set neural encoders?
  • How do sparse storage SpG and parallelism in SpJoin operator perform and benefit the overall performance of SUREL+?
@@ 4.1. Experiment Setup
Extensive experiments have been performed to evaluate SUREL+ using nine homogeneous, heterogeneous, and higher-order homogeneous graphs on three types of tasks: link prediction, relation type prediction, and higher-order pattern prediction. A homogeneous graph is a graph that does not contain node/link types, while a heterogeneous graph includes various node/link types. In our setting, higher-order graphs are hypergraphs consisting of hyperedges connecting two or more nodes.
Datasets Table 2 summarizes the statistics of datasets used to benchmark SGRL methods. Five datasets are selected from the Open Graph Benchmark (OGB, (Hu et al., 2020)) for link and relation type prediction, including social networks of citation - citation2 and collaboration - collab; biological network of protein interaction - ppa and blood vessels - vessel; and one heterogeneous academic network ogb-mag, which contains node types of paper (P), author (A) and their extracted relations. The vessel dataset is a large ($>$3M nodes), sparse, biological graph recently constructed from mouse brains (Paetzold et al., 2021), and has unique significance for examining GRL in scientific discovery. The structure of vessels illustrates the spatial organization of the brain’s microvasculature, which can be used for early detection of neurological disorders, e.g. Alzheimer’s and stroke. Two hypergraph datasets collected by (Benson et al., 2018) are used for higher-order pattern prediction: DBLP-coauthor is a temporal hypergraph, where each hyperedge denotes a time-stamped paper connecting all its authors. tags-math contains groups of tags applied to questions on the website math.stackexchange.com as hyperedges. For higher-order pattern prediction tasks, the number of hyperedges is the main computation bottleneck, in which one may connect more than two nodes. Two industry-level graphs, criteo-click with 16.5M records of online banner ads clicking (Diemert et al., 2017) and twitter-2010 with 1.5B user following relations (Kwak et al., 2010) are used to examine the model scalability for real-world applications.
@@
Baselines We consider two classes of baselines. Canonical GNNs: GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017a), GraphSAINT (Zeng et al., 2020) and their variants with the prefix ‘H*’ that are directly applied for heterogeneous graphs with node types and for hypergraphs through clique expansion. R-GCN (Schlichtkrull et al., 2018) performs relational message passing on heterogeneous graphs. SGRL Models: SEAL (Zhang and Chen, 2018; Zhang et al., 2021), GDGNN (Kong et al., 2022), and SUREL (Yin et al., 2022). SEAL adopts online subgraph sampling due to its intractable space needs for offline extraction. Fig. 3 (a) compares the time cost for subgraph sampling across different SGRL methods. We use all baselines’ official implementations with tuned hyperparameters to match their reported results.
Hyperparameters By default, SUREL+ uses the walk-based sampler, the structural encoder LP, and the better set neural encoder tuned between mean pooling and attention. SUREL+ adopts a 2-layer MLP as $enc(\cdot)$ in Eq. (2) followed by a 2-layer MLP classifier to map set-aggregated readouts for final predictions. Default training hyperparameters: learning rate lr=1e-3 with early stopping of 5 epochs, dropout p=0.1, Adam (Kingma and Ba, 2015) as the optimizer. Analysis of parameters $M$ and $m$ to control the walk-based sampler and $K$ to control the metric-based sampler and selection of structure encoders and set neural encoders are studied in Sec. 4.4.
@@ 4.2. Prediction Accuracy Comparison
Table 3 shows the prediction performance of different methods. SGRL models significantly outperform canonical GNNs on these six link prediction benchmarks, especially on two challenging biological datasets ppa and vessel. Predicting links in biological datasets requires richer structural information that canonical GNNs have limited expressive power to capture. Within SGRL models, SUREL+ achieves comparable performance to SUREL and outperforms SEAL, which validates the effectiveness of the proposed set-based representation for subgraphs. For predictions of relation type and higher-order pattern, we observe additional performance gains (+2$\sim$13%) from SUREL+ compared to SUREL on three of the four datasets. A large performance gap exists between canonical GNNs and SGRL models, particularly in the higher-order case. This demonstrates the inherent limitations of canonical GNNs to make predictions of complex relations involving multiple nodes.
@@ 4.3. Efficiency and Scalability Analysis
Table 4 compares model runtime and memory usage on the four largest benchmarks. SUREL+ offers a reasonable training time compared with canonical GNNs. It shows clear improvement in inference compared to the current SOTA framework SUREL (3-11$\times$ speedups across all datasets) and its predecessor SEAL ($\sim$20$\times$ speedups). SUREL+ achieves comparable and even lower RAM usage than canonical GNNs. Compared to other SGRL models, it can save up to half of RAM with lower usage of GPU SDRAM. This is attributed to set-based subgraphs eliminating node duplicates with structural features, which is further echoed by the analysis in Table 1 and the empirical results in Table 4. The key factor scales SUREL+ to billion-size graphs is its set-based subgraph with the sparse design, while GCN (full adjacency matrix), SEAL (complex subgraph extraction), and SUREL (dense walks with duplicate nodes) are all out of memory (OOM) on twitter-2010.
@@
Fig. 5(a) reports the time cost of different samplers with multithreading on citation2. Fig. 5(b) shows memory consumption to store different types of sampled data (walks in SUREL (Yin et al., 2022) or sets in SUREL+) and associated structural features (LPs, SPDs, PPR scores). Compared to the SUREL sampler, the walk-based sampler in SUREL+ is more efficient and only adds one extra minute for encoding and converting data to SpG format (slash/dash marked in Fig. 5(a)), while achieving 6.94×6.94$\times$6.94 ×, 3.63×3.63$\times$3.63 × and 4.12×4.12$\times$4.12 × memory savings on three OGB datasets, respectively. Those savings are crucial for model scalability as they reduce data transfer from CPU to GPU and reduce GPU operations on duplicate nodes. These two factors dominate the online stage and thus lead to improved memory usage and runtime of SUREL+ in Table 4. In addition, the PPR-based sampler has better scaling performance with more threads. When PPR scores or SPDs are used as structural features, SUREL+ further reduces the memory footprint, though they often slightly harm prediction performance.
Note that, in the above comparison of memory cost, techniques of compressing structural features are adopted both in SUREL (locally) and SUREL+ (globally). When LPs are used as structural features, the two-level indexing in SpG achieves compression of 493×493$\times$493 ×, 11318×11318$\times$11318 ×, 19527×19527$\times$19527 × on three datasets listed in Fig. 5(b).
@@
Fig. 7 shows the speedups and throughput of the SpJoin operator for constructing query-level structural features via multithreading, where the walk join operation of SUREL is used for comparison. SUREL employs a hash-based search for joining walks, which has unfavorable memory access patterns and suffers from imbalanced workloads due to inconsistent searching times across different threads. SUREL+ gains more benefits from multithreading, thanks to sparse arithmetic operations and batch-wise load balancing used in SpJoin.
@@
Figure 6. Comparison of Runtime, Memory Consumption across Different Offline Processing Strategies (the walk-based sampler: $m=4,M=200$, the metric-based sampler: $K=150$). The highlighted areas break down the total consumption w.r.t. (a) sampling, structure encoding, sparse object construction; (b) structural features, node indices/pointers, and sampled walks (SUREL sampler only).
@@ 4.4. Comparison between Different Set Samplers, Structural Features and Set Neural Encoders
Table 5 shows the prediction performance and inference runtime by adopting different combinations of structure encoders and set neural encoders. Landing probabilities (LPs) as structural features perform the best on all three OGB datasets while being the slowest for inference. By recording the landing probabilities over different steps of walks, LPs provide structural information in finer granularity than scalar values of SPDs and PPR scores. Furthermore, the adopted link prediction task might favor more local information held by LPs and SPDs than global information carried by PPR scores. The authors conjecture that other tasks that rely on more global information may favor PPR scores. In comparison, no set neural encoder is always a winner. Attention seems to perform the best on average while slower than mean pooling. LSTM is the slowest. On the two social networks (citation2 and collab), mean pooling can provide comparable prediction results with much fewer parameters. However, prediction on the biological network (ppa) requires more expressive and complicated encoders, where LSTM and attention are favored as they can model more complex interactions between sampled nodes in the union set $\mathcal{S}_{Q}$.
Fig. 8 compares prediction results and inference time by using different hyperparameters $m,M$, and $K$ of set samplers, which heavily affects the coverage of sampled neighborhoods and computation overhead. The performance consistently increases if the walk-based sampler uses a larger $M$, but is not guaranteed for a larger $m$ (broader exploration). Better coverage with a larger $K$ is usually beneficial for the metric-based sampler over citation2 but not for collab, which is due to different characteristics of these two datasets and is also observed by (Yin et al., 2022). In general, small sampling parameters $m~{}(2\sim 4),M~{}(100\sim 400)$ and $K~{}(50\sim 200)$ can yield satisfactory performance with fast inference speed that achieves the trade-off between accuracy and efficiency.
Table 5. Prediction Performance and Inference Time of SUREL+ with Different Combinations of Structure Features (LP, SPD, PPR) and Set Neural Encoders (Mean, LSTM, Attn.). The best and the second best are highlighted in bold and underlined accordingly.
@@ 5. Conclusion
This work proposes a novel framework SUREL+ for scalable subgraph-based graph representation learning. SUREL+ avoids costly subgraph extraction by decoupling it into sampled node sets with structural features, whose join can function as query-induced subgraphs for prediction. SUREL+ benefits from the reusability and compactness of pre-sampled node sets across different queries. Compared to the SOTA framework SUREL, the set-based subgraph of SUREL+ substantially reduces space and time complexity by avoiding heavy node duplication in sampled walks. To handle irregularly sized node sets, SUREL+ designs a customized sparse storage SpG and a sparse join operator SpJoin, providing memory-efficient storage with fast access. In addition, SUREL+ adopts a modular design, enabling users to choose different set samplers, structure encoders, and set neural encoders flexibly based on the nature of their SGRL tasks. Extensive experiments on three types of prediction tasks over nine real-world graph benchmarks show that SUREL+ significantly improves scalability, memory efficiency, and prediction accuracy compared to current SGRL methods and canonical GNNs.
@@ Appendix B More Details
B.1. Other Related Works
Scalable GNN Design. GNNs are the most widely used toolbox for graph representation learning nowadays, although they face certain challenges when directly applied to subgraph-based methods. To address the scalability of GNNs, current studies focus on improving graph subsampling and mini-batch training techniques (Chiang et al., 2019; Zeng et al., 2020). However, graph subsampling used in GNNs fundamentally differs from subgraph extractions in SGRL. The goal of subsampling is to handle GPU memory overflow during full-batch training of GNN models. For SGRL, subgraphs sampled around a query serve as features for making predictions. Consequently, the scaling techniques developed for GNNs cannot be directly applied to SGRL. Another direction is to deploy distributed GNN systems for industry-level graphs. Unfortunately, these specialized techniques, including pipelining (Wan et al., 2022b), partitioned parallelism (Wan et al., 2022a), and update with staleness (Peng et al., 2022) do not address the main bottleneck of subgraph extraction for SGRL methods.
B.2. Model Design
Benefits of Subgraph-based Graph Representation Learning First, subgraph-based representation is versatile for different types of tasks, especially when queries of certain tasks involving multiple nodes and relations, e.g. existence of a link, property of a motif, development of higher-order patterns; while canonical GNNs are limited to handle such polyadic dynamics via node-wise representations (Srinivasan and Ribeiro, 2020; Wang and Zhang, 2021). Second, subgraph-based models are more expressive by pairing with structural features to obtain most expressive structural representations (Srinivasan and Ribeiro, 2020; Li et al., 2020; Wang and Zhang, 2021; Bouritsas et al., 2022). However, canonical GNNs cannot capture intra-distance information and joint relations over multiple nodes, which are critical to distinguishing nodes in structural symmetry and making predictions over them (also refers to the example in Fig. 2). Lastly, subgraph-based methods decouple the model depth from the receptive field since extracted subgraphs are localized to certain hops: when adding more layers for non-linearity, it does not contaminate embedding with irrelevant nodes or get over-smooth as canonical GNNs do. This results in a more robust representation and is particularly beneficial for modeling relations beyond singleton.